# Diffusion in Zero-Shot Learning for Environmental Audio

### v1

@@ 8 Vehicle Engine 17 Speaking 26 Squirrel
TABLE III: Classes in each partition. For ESC-50, random partitions are used from previous work for consistency [18]. FSC22 has not been benchmarked for zero-shot learning before, so a new partition strategy is created with a random number generator.
@@
A modified YAMNet model [19] is trained using only the seen classes in FSC22. A learning rate of $10^{-3}$ is used and early stopping occurred at epoch 259 with an accuracy of $0.6833$. This model is used to retrieve the audio embeddings. The word embedding model is Word2Vec pre-trained on GoogleNews [^1] https://code.google.com/archive/p/word2vec/. Each method uses the same embedding vectors for each respective class and data sample.
IV-A Bilinear Compatibility
@@
The bilinear compatibility network has been used to great success in environmental audio [18, 68, 19]. It represents the state-of-the-art method in zero-shot learning for environmental audio on ESC-50. Therefore, it is important to include it in this comparison.
@@
The bilinear compatibility method has few hyperparameters and the parameters used by Xie and Virtanen [18] for ESC-50 were repeated in our work. FSC22 was tuned to train for 30 epochs.
IV-B CADA-VAE
@@
Cross and distribution aligned variational autoencoder (CADA-VAE) [14] is a successful zero-shot learning method in computer vision, and for this reason it is included in these experiments to give insight into the performance of variational autoencoder methods in environmental audio zero-shot learning.
@@
Hyperparameter tuning was conducted by starting with the default parameters from the computer vision experiments and manually tuning based on the validation sets. Both datasets used the same hyperparameters, aside from the number of classifier epochs. Overall, hyperparameter tuning was more useful in stabilising training and reducing the standard deviation than improving the accuracy. It was found that the large amount of hyperparameters made this method difficult and time consuming to tune, and the original parameters largely did not transfer to ESC-50 and FSC22.
@@
The VAE is trained with warmup strategies for the Kullback–Leibler divergence (KLD), the cross-reconstruction error, and the distance error. The KLD has a factor of one with a warmup over epochs 0 to 93, the cross-reconstruction has a factor of 2.37 with a warmup over epochs 21 to 75, and the distance error has a factor of 5.13 with a warmup over epochs 6 to 22. The VAE’s learning rate is set to $1.5\cdot 10^{-4}$, the batch size is set to 52, and the model is trained for 80 epochs.. The classifier has a learning rate of $10^{-3}$, a batch size of 32 and is trained for 23 epochs for ESC-50 and 20 epochs for FSC22. The loss of the VAE is mean absolute error and the loss of the classifier is negative log likelihood loss. Both use an Adam optimiser. The size of the latent space is 64.
IV-C LisGAN
@@
Leveraging invariant side generative adversarial network (LisGAN) [15] is a more recent GAN method that reports comparatively high results amongst GANs in zero-shot learning. While there are some methodological concerns, these were modified to adhere to the zero-shot learning problem formulation and the core idea of the method remained. A conditional Wasserstein GAN is the main GAN method found in zero-shot learning. However, despite efforts to improve the method, it is still notoriously difficult to train [71, 65]. Smaller experiments were performed with a basic conditional GAN with no success, and often results were no better than random with a high variance across runs. LisGAN showed to be significantly more stable and competitive amongst other zero-shot learning methods.
@@
Hyperparameter tuning of LisGAN involved simple manual tuning, with only a small amount of parameters to tune. These mostly were modified to suit the smaller ESC-50 and FSC22 datasets, as the datasets in computer vision are significantly larger. Both datasets used the same hyperparameters.
@@
LisGAN is used without retraining on unseen classes. The GAN is trained with a batch size of 64, 128 units in the hidden layer of the discriminator and generator, 50 epochs and a learning rate of $10^{-4}$. The classifier is trained with a learning rate of $10^{-2}$ and 50 epochs. The generator, discriminator and classifier use Adam optimisers. The classifier uses a negative log likelihood loss.
ESC-50 FSC22
Val Fold 0 Val Fold 1 Val Fold 2 Val Fold 3 Test Val Test
Bilinear $0.2460\pm 0.0074$ $0.5003\pm 0.0074$ $0.3345\pm 0.0105$ $0.2868\pm 0.0076$ $0.2257\pm 0.0105$ $0.3526\pm 0.0069$ $0.3343\pm 0.0169$
LisGAN $0.2605\pm 0.0227$ $0.4039\pm 0.0220$ $0.2582\pm 0.0187$ $0.2150\pm 0.0195$ $0.2448\pm 0.0165$ $0.299\pm 0.0219$ $0.2142\pm 0.0416$
CADA-VAE $\textbf{0.2915}\pm 0.0180$ $0.3980\pm 0.0181$ $0.2500\pm 0.0251$ $0.3013\pm 0.0251$ $0.2095\pm 0.0337$ $0.3179\pm 0.0137$ $0.3290\pm 0.0339$
ZeroDiffusion $0.2846\pm 0.0111$ $\textbf{0.5279}\pm 0.0283$ $\textbf{0.3593}\pm 0.0189$ $\textbf{0.3342}\pm 0.0224$ $\textbf{0.3091}\pm 0.0278$ $\textbf{0.3975}\pm 0.0132$ $\textbf{0.3467}\pm 0.0200$
TABLE IV: Average accuracy and standard deviation over ten randomly seeded runs for the traditional bilinear compatibility method and the three generative methods for ESC-50 and FSC22. Our method, ZeroDiffusion, outperforms the other methods in accuracy across the partitions of both datasets. The bilinear compatibility method has less variance than the three generative-based methods.
IV-D ZeroDiffusion
@@
ZeroDiffusion is the novel diffusion-based zero-shot learning method introduced in this paper. These experiments aim to determine the performance of this new method in the context of environmental audio. The existing success of diffusion in other areas of machine learning research motivate the creation and experimentation of this method.
The diffusion model is trained using the Adam optimiser with a learning rate of $10^{-3}$ and weight decay of $10^{-4}$, a batch size of 64 and 2000 epochs. Fifty samples are generated from the diffusion model per unseen class. The bilinear compatibility network is trained with a learning rate of $10^{-4}$, batch size of 52, and 20 epochs. These hyperparameters were chosen using manual tuning.
@@ V Results
The results from experiments with the four methods - bilinear compatibility, LisGAN, CADA-VAE, and ZeroDiffusion - using ESC-50 and FSC22 are presented in Table IV.
ZeroDiffusion increases the accuracy on the ESC-50 and FSC22 test partition by $36.95\%$ and $3.71\%$, respectively, from the baseline bilinear method. With the exception of the first validation fold in ESC-50, ZeroDiffusion has a higher accuracy than the other three zero-shot learning methods. This shows that ZeroDiffusion is a new state-of-the-art method for both ESC-50 and FSC22.
Notably, ZeroDiffusion incorporates the bilinear model in its methodology. The difference between the methods is that ZeroDiffusion trains a diffusion model to generate synthetic unseen data and adds these samples to the training of the bilinear method. Because the diffusion model is the only addition, it is reasonable to conclude that the diffusion model is responsible for the improved accuracy.
CADA-VAE has positive performance but is inconsistent across partition strategies. LisGAN has similar inconsistency issues but overall appears to have worse performance across partitions and datasets. This inconsistency continues when considering the standard deviation for each experiment. The generative methods have high variance compared to the baseline method. In the case of ZeroDiffusion, the higher average accuracy and variance are tradeoffs to consider.
@@
The accuracy improvement of ZeroDiffusion supports the use of diffusion models in the zero-shot learning field. Adjacent works with Stable Diffusion [57, 58] are supported by ZeroDiffusion’s performance, indicating that diffusion methods more broadly have potential in a range of domains for zero-shot learning.
Confusion matrices give more insight into model performance through the analysis of class predictions. The confusion matrices for one run of ESC-50 fold 0 and 1 are shown in Tables V and VI, respectively.
In fold 0 of the ESC-50 dataset, the helicopter and mouse click classes have the highest accuracy at 80% and 62.5%, respectively. The helicopter class is predicted at least 70% of the time for four different classes, resulting in a hub around the helicopter class. A hub is a class where predictions across classes are clustered, i.e. the model is much more likely to predict the hubs than other classes. In zero-shot learning, the hubness problem has been identified as a major issue preventing classes from being predicted, resulting in a low accuracy. In fold 0, smaller hubs exist around drinking, brushing teeth and mouse clicking. Other classes are not predicted often. The difference between the hub classes and other classes does not appear to have a clear intuitive explanation. This information could lead to avenues for improvement.
Fold 1 of ESC-50 has a significantly higher accuracy than fold 0. The highest class accuracy is siren at 80%, however all but the clapping class have at least 25% accuracy. Clapping is predicted a total of three times across all clips in this fold, and all three are misclassified hand saw clips. Other classes are predicted more consistently, indicating that the hubness problem isn’t as prevalent in the training of this partition of the dataset.
The confusion matrix in fold 1 is sparse, where classes are only misclassified as a small number of other classes. This phenomenon is also found in traditional supervised learning, where the target class is misclassified as classes with similar features. In zero-shot learning, it is the machine learning-based semantic embeddings that are being compared, and so intuition from traditional supervised learning may not apply.
Some misclassifications appear intuitive, such as the case where the sound of a washing machine is incorrectly labeled as a helicopter. Both classes are loud and mechanical, motivating a potential misunderstanding from the model. Fireworks are misclassified as mouse clicking in fold 0, and the reason may not initially be intuitive, but both sounds are comprised of short, sharp noises. There is a difference in loudness, but the audio data is normalised before input to the network and so this difference may be reduced. While there are some intuitive misclassifications, there are also non-intuitive misclassifications. Pig is incorrectly labeled as brushing teeth, and door knocking is mislabeled as rain.
To gain a deeper understanding, dimensionality reduction with t-distributed stochastic neighbour embedding (t-SNE) [72] on the audio embeddings produces Figure 3 for fold 0 and Figure 4 for fold 1. In fold 0, pig (class 2, dark purple) is spread across a wide area, including overlapping with brushing teeth (class 27, dark blue). This relationship gives insight into the non-intuitive misclassification of pig.
Fold 1 has a significantly higher accuracy than fold 0, as shown in Table IV. The t-SNE graph for fold 1 shows more distinct clusters than fold 0, indicating that this clustering property of the audio embeddings is important in correctly classifying unseen classes. This is reinforced when considering the clapping class, which both is the poorest performing class in fold 1 and has the greatest spread across the t-SNE graph.
The t-SNE graphs reveal insights into the hubness problem, and asymmetry in the incorrect predictions further highlights the problem. Pig is mislabeled as brushing teeth more than 50% of the time, while brushing teeth is never mislabeled as pig. This shows that certain classes are more likely to be predicted than others. The classes in fold 0 that are frequently predicted as the helicopter class are closer together in the t-SNE graph, showing their similarity in the embedding space. This similarity is a potential cause of the hubness problem.
Class index 38 29 35 3 40 2 27 46 31 48
38 1 9 0 3 0 3 9 0 14 1
29 1 18 0 0 1 0 18 0 1 1
35 1 0 6 0 30 1 1 1 0 0
3 0 0 5 0 29 5 1 0 0 0
40 0 0 6 0 32 0 0 1 0 1
2 0 6 3 0 3 5 21 0 0 2
27 2 34 0 1 0 0 1 0 2 0
46 0 0 2 0 28 2 0 8 0 0
31 1 5 0 0 2 0 6 0 25 1
48 4 6 0 0 2 0 8 1 15 4
TABLE V: Confusion Matrix for fold 0 of the ESC-50 dataset. Rows represent the true class, and columns represent the predictions. Many predictions are clustered around class 40, with smaller hubs on classes 29, 27 and 31.
Figure 3: ESC-50 fold 0 audio embeddings with t-SNE dimensionality reduction. Class 40 (helicopter), 46 (church bells), 35 (washing machine) and 3 (cow) are all concentrated to the right side of the scatterplot. The confusion matrix in Table V shows that all of these classes are classified as helicopter a majority of the time.
Class index 39 36 42 13 32 22 19 49 26 21
39 10 0 0 0 9 0 2 3 0 16
36 0 24 3 4 0 0 6 1 0 2
42 0 0 32 3 0 0 0 2 3 0
13 0 4 7 14 1 0 5 0 1 8
32 2 0 0 0 23 0 2 13 0 0
22 14 1 0 1 3 0 14 6 1 0
19 0 5 10 0 0 0 22 0 0 3
49 1 3 0 1 0 3 0 11 8 13
26 4 0 2 4 0 0 0 4 15 11
21 2 0 0 1 1 0 0 0 8 28
TABLE VI: Confusion matrix for fold 1 of the ESC-50 dataset. Rows represent the true class, and columns represent the predictions. There are small hubs on classes 42, 19, 26 and 21 and all classes are at least partially correctly identified, except for class 22 which is never correctly classified.
Figure 4: ESC-50 fold 1 audio embeddings with t-SNE dimensionality reduction. Fold 1 exhibits more defined class clusters than fold 0. Class 22 (clapping, dark blue) has a greater spread, and is the poorest classified class as shown in Table VI.
t-SNE graphs give a deeper insight into issues surrounding zero-shot learning, and can indicate avenues for improvement. The relationship between the clustering properties of the audio embeddings and model accuracy show that embeddings have an effect on the performance of a zero-shot learning model. More defined clustering in the audio embeddings produces higher accuracy, and hubs in the confusion matrix are reflected in the t-SNE diagram. The confusion matrices highlight the abilities of the model beyond an accuracy value. The sparsity of the matrices demonstrate that the model recognises the similarity of clips from the same class, and when it correctly maps these to the correct class the performance is very high.

### v2

@@ IV-A6 GTZAN
ARCA23K-FSD fold 0 crash cymbal, run, zipper (clothing), acoustic guitar, gong, knock, train, crack, cough, cricket
fold 1 electric guitar, chewing mastication, keys jangling, female speech woman speaking, crumpling crinkling, skateboard, computer keyboard, bass guitar, stream, toilet flush
fold 2 tap, water tap faucet, squeak, snare drum, finger snapping, walk footsteps, meow, rattle (instrument), bowed string instrument, sawing
fold 3 rattle, slam, whoosh swoosh swish, hammer, fart, harp, coin (dropping), printer, boom, giggle
fold 4 clapping, crushing, livestock farm animals working animals, scissors, writing, wind, crackle, tearing, piano, microwave oven
fold 5 trumpet, wind instrument woodwind instrument, child speech kid speaking, drill, thump thud, drawer open close, male speech man speaking, gunshot gunfire, burping eructation, splash splatter
fold 6 female singing, wind chime, dishes pots pans, scratching (performance technique), crying sobbing, waves surf, screaming, bark, camera, organ
@@
TABLE I: Classes in each partition. For ESC-50, random partitions are used from previous work for consistency [19]. All other datasets have not been benchmarked for zero-shot learning before, so new partitions are created using a random number generator.
IV-B Embeddings
A modified YAMNet model [20] - a YAMNet architecture with layers removed and an embedding layer added in before the output layer - is trained using only the seen classes in each dataset partition. For all datasets, the embedding layer size is 128, the learning rate is $10^{-3}$, and the batch size is 16. The models are trained with early stopping and a patience of 200. This model is then used to retrieve audio embeddings to train each zero-shot learning method.
@@ IV-B Embeddings
The word embedding model used is Word2Vec pre-trained on GoogleNews [^1] https://code.google.com/archive/p/word2vec/. Each method uses the same embedding vectors for each respective class and data sample.
IV-C Non-linear ALE
@@
The non-linear ALE compatibility network has been used to great success in environmental audio [48, 20]. It represents the state-of-the-art method in zero-shot learning for environmental audio on ESC-50. Therefore, it is important to include it in this comparison.
@@
The non-linear compatibility method has few hyperparameters and the parameters used by Xie et al. [48] for ESC-50 were repeated in our work for all datasets, except FSC22 which was tuned to train for 30 epochs.
IV-D CADA-VAE
@@
Cross and distribution aligned variational autoencoder (CADA-VAE) [15] is a successful zero-shot learning method in computer vision. For this reason, it is included in these experiments to give insight into the performance of variational autoencoder methods in environmental audio zero-shot learning.
@@
Hyperparameter tuning was conducted by starting with the default parameters from the computer vision experiments and manually tuning based on the validation sets. Both datasets used the same hyperparameters, aside from the number of classifier epochs. Overall, hyperparameter tuning was more useful in stabilising training and reducing the standard deviation than improving the accuracy. It was found that the large amount of hyperparameters made this method difficult and time-consuming to tune, and the original parameters largely did not transfer to ESC-50 and FSC22.
@@
The final training configuration was determined iteratively, guided by validation performance. The VAE is trained with warmup strategies for the Kullback–Leibler divergence (KLD), the cross-reconstruction error, and the distance error. The KLD has a factor of one with a warmup over epochs 0 to 93, the cross-reconstruction has a factor of 2.37 with a warmup over epochs 21 to 75, and the distance error has a factor of 5.13 with a warmup over epochs 6 to 22. The VAE’s learning rate is set to $1.5\cdot 10^{-4}$, the batch size is set to 52, and the model is trained for 80 epochs. The classifier uses a learning rate of $10^{-3}$ and a batch size of 32. It is trained for 23 epochs on ESC-50; 20 epochs on FSC22, UrbanSound8K and GTZAN; 30 epochs on TAU 2019; and 50 epochs on ARCA23K-FSD, with the number of epochs selected based on the size of each dataset. The loss of the VAE is a mean absolute error and the loss of the classifier is a negative log-likelihood loss. Both use an Adam optimiser. The size of the latent space is 64.
IV-E LisGAN
@@
Leveraging invariant side generative adversarial network (LisGAN) [16] is a more recent GAN method that reports comparatively high results amongst GANs in zero-shot learning. LisGAN is a transductive method while we consider only the inductive case in these experiments. The transductive component is removed and the core idea of the method remains. A conditional Wasserstein GAN [38] is the main GAN method found in zero-shot learning. However, despite efforts to improve the method, it is still notoriously difficult to train [57, 43]. Smaller experiments were performed with a basic conditional GAN with no success, and often results were no better than random with a high variance across runs. LisGAN has shown to be significantly more stable and competitive amongst other zero-shot learning methods, even with the transductive component removed.
@@
Hyperparameter tuning of LisGAN involved simple manual tuning, with only a small amount of parameters to tune. These were modified to suit the smaller audio datasets, as the datasets in computer vision are significantly larger. All datasets used the same hyperparameters.
@@
LisGAN is used without retraining on unseen classes. The GAN is trained with a batch size of 64, 128 units in the hidden layer of the discriminator and generator, 50 epochs and a learning rate of $10^{-4}$. The classifier is trained with a learning rate of $10^{-2}$ and 50 epochs. The generator, discriminator and classifier use Adam optimisers. The classifier uses a negative log-likelihood loss.
ALE LisGAN CADA-VAE ZeroDiffusion
ESC-50 fold 0 $0.2278\pm 0.0091$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2605}}\pm 0.0227$ $\textbf{0.2915}\pm 0.0180$ $0.2598\pm 0.0140$
fold 1 $\textbf{0.4932}\pm 0.0093$ $0.4039\pm 0.0220$ $0.3980\pm 0.0181$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.4665}}\pm 0.0194$
fold 2 $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2937}}\pm 0.0129$ $0.2582\pm 0.0187$ $0.2500\pm 0.0251$ $\textbf{0.3110}\pm 0.0082$
fold 3 $0.2680\pm 0.0117$ $0.2150\pm 0.0195$ $\textbf{0.3013}\pm 0.0251$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2975}}\pm 0.0127$
test $0.2340\pm 0.0167$ $\textbf{0.2448}\pm 0.0165$ $0.2095\pm 0.0337$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2343}}\pm 0.0118$
average $0.3033\pm 0.0119$ $0.2765\pm 0.0199$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.3102}}\pm 0.0240$ $\textbf{0.3138}\pm 0.0132$
ARCA23K-FSD fold 0 $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2420}}\pm 0.0032$ $0.2183\pm 0.0128$ $0.2398\pm 0.0125$ $\textbf{0.2879}\pm 0.0153$
fold 1 $0.2876\pm 0.0028$ $0.3059\pm 0.0107$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.3142}}\pm 0.0149$ $\textbf{0.3418}\pm 0.0150$
fold 2 $\textbf{{\color[rgb]{0.5,0.5,0.5}0.3723}}\pm 0.0061$ $0.3086\pm 0.0121$ $0.3164\pm 0.0120$ $\textbf{0.4819}\pm 0.0188$
fold 3 $\textbf{0.2546}\pm 0.0029$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2113}}\pm 0.0148$ $0.2036\pm 0.0101$ $0.1885\pm 0.0200$
fold 4 $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2283}}\pm 0.0020$ $0.2173\pm 0.0113$ $0.2126\pm 0.0140$ $\textbf{0.2833}\pm 0.0279$
fold 5 $0.3205\pm 0.0050$ $0.3196\pm 0.0142$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.3567}}\pm 0.0144$ $\textbf{0.4335}\pm 0.0321$
test $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2129}}\pm 0.0051$ $0.1772\pm 0.0068$ $\textbf{0.2414}\pm 0.0145$ $0.2081\pm 0.0261$
FSC22 val $0.2650\pm 0.0088$ $0.2990\pm 0.0219$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.3179}}\pm 0.0137$ $\textbf{0.3269}\pm 0.0113$
test $\textbf{{\color[rgb]{0.5,0.5,0.5}0.2968}}\pm 0.0070$ $0.2142\pm 0.0416$ $\textbf{0.3290}\pm 0.0339$ $0.2775\pm 0.0071$
UrbanSound8k $\textbf{0.4997}\pm 0.0071$ $0.3848\pm 0.0238$ $0.4741\pm 0.0808$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.4961}}\pm 0.0777$
TAU2019 $\textbf{{\color[rgb]{0.5,0.5,0.5}0.4377}}\pm 0.004$ $0.3878\pm 0.0188$ $0.3158\pm 0.0248$ $\textbf{0.4857}\pm 0.0395$
GTZAN $0.5705\pm 0.0109$ $\textbf{0.5824}\pm 0.0287$ $0.5650\pm 0.0463$ $\textbf{{\color[rgb]{0.5,0.5,0.5}0.5803}}\pm 0.0281$
Average $\textbf{{\color[rgb]{0.5,0.5,0.5}0.3227}}\pm 0.0076$ $0.2936\pm 0.0187$ $0.3137\pm 0.0242$ $\textbf{0.3486}\pm 0.0221$
TABLE II: Average accuracy and standard deviation over ten randomly seeded runs for the traditional non-linear ALE method and the three generative methods for the six chosen datasets. Our method, ZeroDiffusion, is similar or superior in almost all cases, with the exception of ESC-50 where no one method consistently performs the best. When averaging ESC-50 results, ZeroDiffusion is ahead of the other methods. The best accuracy in each row is bolded and the second-best accuracy is bolded in grey. The average value across all datasets is displayed in the final row, where ZeroDiffusion has the highest average accuracy and ALE reports the lowest overall variance.
IV-F ZeroDiffusion
@@
ZeroDiffusion is the novel diffusion-based zero-shot learning method introduced in Section III-C. These experiments aim to determine the performance of this new method in the context of environmental audio. The existing success of diffusion in other areas of machine learning research motivates the creation and experimentation of this method.
The diffusion model is trained using the Adam optimiser with a learning rate of $10^{-3}$ and weight decay of $10^{-4}$, a batch size of 64 and 50 epochs. The number of samples generated for each unseen class corresponds to the average number of samples per class in the dataset. For TAU 2019, this is 1,440 samples per class; UrbanSound8k is unbalanced but averages 873 samples per class; GTZAN is 100 samples per class; ESC-50 uses 40 samples per class; and finally, ARCA23K-FSD is unbalanced and averages 339 per class. We train the non-linear compatibility network with a learning rate of $10^{-4}$, batch size of 64, and 10 epochs. For the diffusion loss terms, reconstruction loss was weighted $1.0$, maximum mean discrepancy at $1.0$, variance loss as $0.1$, mean reconstruction loss at $0.2$, and cosine similarity at $2.0$.
These hyperparameters were chosen using manual tuning across all validation sets. The same parameters were used for all experiments except for the synthetic dataset size, which changed relative to the dataset size.
@@ V Results
Table II presents the results from experiments with the four methods - non-linear ALE, LisGAN, CADA-VAE, and ZeroDiffusion - using ESC-50, ARCA23K-FSD, FSC22, UrbanSound8k, TAU 2019 and GTZAN.
ZeroDiffusion maintains an equally top result for UrbanSound8k and GTZAN, shared with ALE and LisGAN, while it outperforms other methods on TAU 2019. Notably, despite not outperforming all methods across all three smaller datasets, it is the only method to maintain a top result across all three. For the FSC22 dataset, the results are worse for ZeroDiffusion but it outperforms other methods on the validation set and underperforms on the test set. This indicates an issue with reliability for the new method; however, it is clear that the other presented methods also lack reliability.
@@
Overall, the accuracy improvement of ZeroDiffusion across most datasets and partitions supports the use of diffusion models in the field of zero-shot learning. Adjacent works with Stable Diffusion [44, 45] are supported by ZeroDiffusion’s performance, indicating that diffusion methods more broadly have potential in a range of domains for zero-shot learning. ZeroDiffusion is agnostic to the specific data type and works entirely with embeddings, similarly to ALE, and therefore is a more flexible method than the image domain-specific Stable Diffusion methods.
V-A Comparison of Dataset Sizes
ARCA23K-FSD is a significantly larger dataset than any of the other datasets, and as such allows for a deeper 7-fold analysis. ZeroDiffusion outperforms the other methods in almost every case, and in the cases of folds 0, 2, 4 and 5 it has a significant improvement over any other method. This result indicates that ZeroDiffusion may draw benefits from the larger dataset, with richer embeddings produced by the diffusion model as a result of more classes available. However, despite the overall success, fold 3 extremely underperforms and could provide an avenue for improvement.
Whilst ESC-50 is not as large as ARCA23K-FSD, its large class size allows for 5-fold partitioning. However, the results for this dataset are extremely unstable when considering the best-performing methods. No one method consistently outperforms this dataset, and all methods outperform the others on at least one partition. To gain a better understanding, the average across all folds is calculated and ZeroDiffusion marginally has the highest accuracy. An indication of why this is happening may be revealed in the average number of audio clips per class in each dataset. ESC-50 has the lowest with 40 samples per class. This may directly result in the method failing to generalise. FSC22 is the next largest dataset per class with 75 samples, and this dataset similarly experienced issues with performance. Of the smaller datasets, TAU 2019 has 1,440 per class, the highest of any dataset in the experiments, and ZeroDiffusion produced a significant improvement over other methods. ARCA23K-FSD has on average 339 samples per class, a reasonable increase over both ESC-50 and FSC22, and reports good results. Given this is an unbalanced dataset, this may provide insight into why ZeroDiffusion underperforms on some folds in ARCA23K-FSD.
V-B Ablation
Notably, ZeroDiffusion incorporates the non-linear model in its methodology. The ALE results therefore represent an ablation of ZeroDiffusion without the diffusion-generated data component. The only other changes present are the weight decay term on the classifier ($10^{-4}$ for ALE and $10^{-5}$ for ZeroDiffusion) and the loss function (WARP for ALE and cross-entropy for ZeroDiffusion’s classifier); however, tests were done to align these with poorer results on both ALE and ZeroDiffusion across datasets when changing either or both terms. Given all of this, is it reasonable to conclude that the addition of a diffusion model to generate synthetic data for classifier training is responsible for the improved accuracy between ALE and ZeroDiffusion.
V-C Variance
Standard deviation is recorded for each experiment, and it is clear that across runs the generative-based methods have a high standard deviation. Meanwhile, ALE has a relatively low standard deviation. This indicates that generative-based methods, including ZeroDiffusion, are more variable and potentially unreliable compared to the more stable ALE method. Additionally, generative methods typically have a wide range of hyperparameters. Methods such as ALE have a low number of hyperparameters, making them user-friendly and simple to apply to new problems.